# Further author information: (Send correspondence to E.B.)

### v3

@@ 3.3 CrossViewFusion for Unified Multi-View Integration
[…]
@@ 4.2 Implementation Details
Videos are center-cropped to $\times 224$ 224, rescaled to the [0,1][0,1][ 0 , 1 ] range, and normalized with mean 0.450.450.45 and standard deviation 0.2250.2250.225. Frames are sampled uniformly with temporal interpolation of positional embeddings for variable lengths.
@@ 5.3 Per-Domain Accuracy Analysis
Table 3 reports per-scenario accuracy across all models and configurations. SkillFormer consistently outperforms baselines in the Ego+Exos setting for structured and physically grounded activities such as Basketball (77.88%), Cooking (60.53%), and Bouldering (31.87%). These domains benefit from synchronized egocentric and exocentric perspectives, which enable better modeling of spatial layouts and temporally extended actions. The fusion of multi-view signals allows SkillFormer to exploit cross-perspective cues, such as object-hand interactions or full-body movement trajectories, which are often ambiguous or occluded in single-view inputs.
Interestingly, in Music, the Ego-only configuration achieves the highest accuracy (72.41%), suggesting that head-mounted views are sufficient to capture detailed instrument manipulation, while additional views may introduce redundant or misaligned information. This highlights the flexibility of SkillFormer to adapt its fusion mechanism to view-specific signal quality.

### v4

@@ 3.3 CrossViewFusion for Unified Multi-View Integration
$\mathbf{X}_{attn},$\_ $=\text{MultiheadAttention}(Q=\mathbf{X}_{norm},K=\mathbf{X}_{norm},V=\mathbf{X}_{norm})$ (2)
@@ 4.2 Implementation Details
Videos are center-cropped to $224\times 224$, rescaled to the $[0,1]$ range, and normalized with mean $0.45$ and standard deviation $0.225$. Frames are sampled uniformly with temporal interpolation of positional embeddings for variable lengths.
@@ 5.3 Per-Domain Accuracy Analysis
Table 3 reports per-scenario accuracy across all models and configurations. SkillFormer consistently outperforms baselines in the Ego+Exos setting for structured and physically grounded activities such as Basketball (77.88%) and Cooking (60.53%). These domains benefit from synchronized egocentric and exocentric perspectives, which enable better modeling of spatial layouts and temporally extended actions. The fusion of multi-view signals allows SkillFormer to exploit cross-perspective cues, such as object-hand interactions or full-body movement trajectories, which are often ambiguous or occluded in single-view inputs.
Interestingly, in Music, the Ego-only configuration achieves the highest accuracy (72.41%), suggesting that head-mounted views are sufficient to capture detailed instrument manipulation, while additional views may introduce redundant or misaligned information. Similarly, in Bouldering, the Exos-only configuration outperforms both Ego (30.77%) and Ego+Exos (31.87%) with 33.52% accuracy, indicating that third-person perspectives better capture full-body spatial positioning and climbing technique assessment. This highlights the flexibility of SkillFormer to adapt its fusion mechanism to view-specific signal quality, but also demonstrates that domain-specific viewpoint advantages can be diluted when combining perspectives that provide conflicting or redundant information.